OpenStudy (anonymous):

will fan and medal!!! Solve the following problem for the roots by using the quadratic formula. 2(6 - x) = x(x + 5)

OpenStudy (luigi0210):

You know what the quadratic formula is right?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Great, so I can assume you know what the problem is with this question huh? It's not in the right form :P So what you need to do in order to solve it with the quadratic formula is distribute.

OpenStudy (anonymous):

okay so it would be 12-2x=x^2+5x @Luigi0210

OpenStudy (luigi0210):

Right, now move everything to one side, any side, so the other side equals zero. Put it in this form: \(\large \color{green}{Ax^2+Bx+C=0}\)

OpenStudy (anonymous):

Okay -x^2-5x+12-2x=0 -x^2-7x+12=0 @Luigi0210

OpenStudy (luigi0210):

Yes, now just multiply everything by a negative to get: \(\large x^2+7x-12=0\) Now we got it in standard form. So all we need to do from here is plug in the values into the quadratic formula: \(\Large \frac{-\color{blue}{b} \pm \sqrt{(\color{blue}{b})^2-4(\color{red}{a})(\color{green}{c)} } } {2(\color{red}{a})} \) \(\Large \color{red}{1}x^2 +\color{blue}{7}x-\color{green}{12} \) There, I color-coded it too :3 This lag is terrible too, sorry about that